I am working on a review and keep missing one problem. It doesn't say which one I am getting wrong. Can you please look over these and tell me what I am missing?

I am going to sum up the questions.

1) Total Assets? I answered $4,140 counting cash and supplies only

Accounts Payable $1394

Cash $3776

Dividend $507

Fuel Expense $477

Interest Expense $393

Notes Payable $2466

Retained Earnings $285

Service Revenue $5489

Supplies $364

2) Total Expenses in Feb? I answered $1,843 using $1194 + $649

a) Paid $1,194 for internet advertisement run in Feb.

b) Paid utility bill $1,247 for Jan services

c) Paid $5,245 wages, $649 for work in Feb and the rest for work in Jan

d) Purchased $991 of supplies in cash

3) Total net increase to assets in Feb? I answered $11,149. I used $8,166 + $2,983

a) Received $9,207 cash from customers for membership services that had been earned and billed in Jan

b) Received $8,166 cash from customers sales earned in Feb.

c) Received $2,983 cash for an advance on a company party held in March. (This is a sports company customers can rent space from)

4) What type of account is "Notes Payable"? I answered liability

5) Sold 2690 subscriptions for $46 each. $16,725 paid in cash and the rest to be paid at a later time. What amount is Accounts Receivable?

2690 * 46 = 123,740 - cash $16,725 = Accounts Receivable $107,015

6) What is the total revenue in Feb? I answered $6,660

a) Received $6,491 cash from customers for services earned and billed in Jan.

b) Received $6,660 cash from customers sales revenue earned in Feb.

c) Received $2,109 cash advance on a company party held in March (same company as above, renting their space to customers)

7) How much is S.E. ? I answered $2,710

Cash $5,378 Notes Payable $5,289 Common Stock $2,710

8) What is the end balance of Accounts Payable? I answered $564

Beginning Balance is $570, $706 in credits, and $712 in debits
Answer #1

Profit = revenue - expenses = 5489 - 477 -393 = 4619

retained earnings = opening bal + profit - dividend = 285 + 4619 - 507 = 4397

total liabilities = retained earnings + accounts payable + notes payable = 4397 +1394 + 2466 = 8257 is total assets

answer 2 to 8 are absolutely correct. The only answer which has some confusion is part 1 of total assets. if the answer still does not match, please comment below.